f(x)=log(base 5)x;
Solve and find
A) x=1
B) x=125

A) f(1)= log(base 5)(1)
= 5log(base5)(1)
= 1
B) f(125)= log(base 5)(125)
= 5^x=125
= log5^x= log125
= xlog5=log125
= xlog5/log5= log125/log5
= 3
Were these two problems done correctly?

Answer by stanbon(75887)

f(x)=log(base 5)x;
Solve and find
A) x=1
f(1) = log5(1) = log(1)/log(5) = 0/5 = 0
f(1) = 0
------
B) x=125
f(125) = log5(125) = log(125)/log(5) = 3
------
Cheers,
Stan H.
=========